Solve the equation of mod for x =|x×x-5x+2|=4 and |x×x-3x+2|=2

Grade:11

1 Answers

Vikas TU
6 years ago
. |x2-5x+2|=4
          x2-5x+2=+/-4
          case 1:
          x2-5x+2=4
          x2-5x-2=0
          x=5+root33/2, 5-root33/2
case 2:
x2-5x+2=-4
x2-5x+6=0
x=3,2
          |x2-3x+2|=2
          x2-3x+2=+/-2
case 1:
          x2-3x+2=2
          x2-3x=0
          x2=3x
          x=3,-3
case 2:
          x2-3x+2=-2
          x2-3x+4=0
          x=3/2+iroot7/2, 3/2-iroot7/2